OpenStudy (anonymous):

what are the foci of (x-3)^2/169+(y+5)^2/144=1 (–8, 5) and (2, 5)

jimthompson5910 (jim_thompson5910):

unfortunately this is incorrect

OpenStudy (anonymous):

ok can you tell me how to do this ?

jimthompson5910 (jim_thompson5910):

the major axis is the longer axis

jimthompson5910 (jim_thompson5910):

sqrt(169) = 13 sqrt(144) = 12 so 13*2 = 26 is the length of the major axis and it is horizontal since it's under the x^2 term

jimthompson5910 (jim_thompson5910):

the center of this ellipse is what point?

OpenStudy (anonymous):

-5,3 ?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

Hint: the center of (x-h)^2/a^2+(y-k)^2/b^2=1 is (h,k)

OpenStudy (anonymous):

so is it -3,5 ? what are the foci of (x{{{-3}}})^2/169+(y{{{+5}}})^2/144=1

jimthompson5910 (jim_thompson5910):

closer, but you have your signs mixed up

jimthompson5910 (jim_thompson5910):

the center is actually (3,-5)

jimthompson5910 (jim_thompson5910):

so (h,k) = (3,-5)

jimthompson5910 (jim_thompson5910):

now we need to find the focal distance c c^2 = a^2 - b^2 c^2 = 13^2 - 12^2 c^2 = 169 - 144 c^2 = 25 c = sqrt(25) c = 5 This is the distance from the center to either focus (and you go along the major axis)

jimthompson5910 (jim_thompson5910):

So one focus is at (h-c, k) Another focus is at (h+c, k)

jimthompson5910 (jim_thompson5910):

h = 3 k = -5 c = 5

OpenStudy (anonymous):

so plug it in ? (3-5, -5)

jimthompson5910 (jim_thompson5910):

yes, so that gives you (-2, -5)

jimthompson5910 (jim_thompson5910):

do the same for (h+c, k)

OpenStudy (anonymous):

(3+5, -5) (8,-5)

jimthompson5910 (jim_thompson5910):

good, your two foci are: (-2,-5) and (8,-5)

OpenStudy (anonymous):

omg thankyouu (:. ive been asking for help all day. and no one was helping . i greatly appreciate it. (:. i have more okk ?

jimthompson5910 (jim_thompson5910):

how much more?

jimthompson5910 (jim_thompson5910):

and i'm sry no one else would help

OpenStudy (anonymous):

like 6. you can just tell me if its right or wrong if you want.

jimthompson5910 (jim_thompson5910):

alright that sounds like a plan

jimthompson5910 (jim_thompson5910):

what you get wrong, try to work them again and come back

OpenStudy (anonymous):

okk. want me to just put them here ?

jimthompson5910 (jim_thompson5910):

either way works

OpenStudy (anonymous):

1)What is the focus of the parabola y = -1/16(x+3)²-6 (–3, –10) 2)given the ellipse ((x - 4)^2)/169 + ((y + 8)^2)/100 = 1 and the hyperbola ((x - 4))^2)/169 + ((y + 8)^2)/100 = 1, which has a greater value? hyperbola 3)What is the transverse axis and conjugate axis of the hyperbola y²/81-x²/121=1 The transverse axis is 121 and the conjugate axis is 81.

jimthompson5910 (jim_thompson5910):

1)What is the focus of the parabola y = -1/16(x+3)²-6 (–3, –10) correct

jimthompson5910 (jim_thompson5910):

#2 doesn't make much sense...what two values are you comparing exactly?

OpenStudy (anonymous):

im not sure what the question was asking , i just guessed. ellipse hyperbola neither cannot be determine

OpenStudy (anonymous):

those were the options,

jimthompson5910 (jim_thompson5910):

is that the full question for #2?

OpenStudy (anonymous):

yess.

jimthompson5910 (jim_thompson5910):

hmm not sure what to make of #2 then #3 is incorrect

OpenStudy (anonymous):

okk . ill try and figure it out again. thankyou (:

jimthompson5910 (jim_thompson5910):

yw